OpenStudy (anonymous):

the table numbers are different I'm still solving for K... X: 20 30 40 Y: 2 10 18

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Same process as before.

OpenStudy (jhannybean):

But is this varying directly or inversely?

OpenStudy (anonymous):

\[y=kx\\ \frac{y}{x}=k\] check to see if \[\frac{2}{20}=\frac{10}{30}=\frac{18}{40}\]

OpenStudy (anonymous):

directly

OpenStudy (jhannybean):

@satellite73 beat em to it.

OpenStudy (jhannybean):

me*

OpenStudy (anonymous):

but y doesn't vary with x right??

OpenStudy (jhannybean):

Why wouldnt it? When x increases, y decreases, and vise versa

OpenStudy (anonymous):

well my answer key says that it doesn't... again I just need help showing work

OpenStudy (anonymous):

the fractions i wrote above are not equal`

OpenStudy (anonymous):

since they are not equal, it does not vary "directly"

OpenStudy (anonymous):

that is the work there is nothing else to do

OpenStudy (anonymous):

Oh I see now thank you